OpenStudy (anonymous):

0.3x-0.2y=4 0.4x+0.5y=- 55/77 the neg is for the whole fraction not just 55 What is the order pair please?

OpenStudy (anonymous):

The ordered paid is: {x -> 8.07453, y -> -7.8882}

OpenStudy (anonymous):

how sure are you on that one?

OpenStudy (anonymous):

Check the results by plugging in the solution for x and y into the original equations using Mathematica. Would you like the expressions posted here?

OpenStudy (anonymous):

It also solve with the elimination method

OpenStudy (anonymous):

Your equations with reduced rationalized coefficients\[\left\{\frac{3 x}{10}-\frac{y}{5}=4,\frac{2 x}{5}+\frac{y}{2}=-\frac{5}{7}\right\} \] Can you proceed with a solution on your own?

OpenStudy (anonymous):

I am not sure exactly how to do that? please help

OpenStudy (anonymous):

working

OpenStudy (anonymous):

Getting rid of the fractional coefficients yeilds: \[\{-40+3 x-2 y=0,50+28 x+35 y=0\} \] The idea is to multiply each side on one or both equations with constants so that when the equations are summed, one of the variables are eliminated and you end up with one equation in one unknown.

OpenStudy (anonymous):

those are huge answer though to plug in to my paper.

OpenStudy (anonymous):

-1400 + 105 x - 70 y == 0 100 + 56 x +70 y == 0

OpenStudy (anonymous):

-1300 + 161 x == 0 x=1300/161

OpenStudy (anonymous):

So is that the order pair 1300,161?

OpenStudy (anonymous):

\[\frac{1300}{161}=8.07453 \] Which is the answers for x that I gave you in the first place. When I supplied you with the original answers I had not examined the numerical coefficients to see if they could be rationalized. I used Mathematica for finding the solutions. An understood rule for this program is that if any number appears with a decimal point, the numbers in the solution will be decimal numbers.

OpenStudy (anonymous):

The final ordered pair is solving with the fractional equation forms is: \[\left\{x\to \frac{1300}{161},y\to -\frac{1270}{161}\right\} \]